# You Don’t Need Pre-built Graphs for RAG: Retrieval Augmented Generation with Adaptive Reasoning Structures

### v1

@@ Abstract
Large language models (LLMs) often suffer from hallucination, generating factually incorrect statements when handling questions beyond their knowledge and perception. Retrieval-augmented generation (RAG) addresses this by retrieving query-relevant contexts from knowledge bases to support LLM reasoning. Recent advances leverage pre-constructed graphs to capture the relational connections among distributed documents, showing remarkable performance in complex tasks. However, existing Graph-based RAG (GraphRAG) methods rely on a costly process to transform the corpus into a graph, introducing overwhelming token cost and update latency. Moreover, real-world queries vary in type and complexity, requiring different logic structures for accurate reasoning. The pre-built graph may not align with these required structures, resulting in ineffective knowledge retrieval. To this end, we propose a Logic-aware Retrieval-Augmented Generation framework (LogicRAG) that dynamically extracts reasoning structures at inference time to guide adaptive retrieval without any pre-built graph. LogicRAG begins by decomposing the input query into a set of subproblems and constructing a directed acyclic graph (DAG) to model the logical dependencies among them. To support coherent multi-step reasoning, LogicRAG then lin- earizes the graph using topological sort, so that subproblems can be addressed in a logically consistent order. Besides, LogicRAG applies graph pruning to reduce redundant retrieval and uses context pruning to filter irrelevant context, significantly reducing the overall token cost. Extensive experiments demonstrate that LogicRAG achieves both superior performance and efficiency compared to state-of-the-art baselines.
@@ Introduction
Large language models (LLMs), like Claude (Anthropic 2024) and ChatGPT (OpenAI 2023), have shown remarkable ability in a wide range of tasks like complex reasoning (Chen et al. 2024c; Bi et al. 2024b), question answering (Tan et al. 2025; Khashabi et al. 2020), content generation (Chowdhery et al. 2023; Zhang et al. 2024a). However, these foundation models frequently struggle with domain-specific tasks, often producing hallucinations or inaccurate responses when handling knowledge-intensive queries (Tan et al. 2024; Huang et al. 2023; Fang et al. 2024, 2025; Bi et al. 2025a, 2024d). To mitigate these issues, retrieval-augmented generation (RAG) (Gao et al. 2023; Lewis et al. 2020; Bi et al. 2024a, 2025b; Zhang et al. 2025a) emerges as a promising framework that enhances LLMs by retrieving query-relevant contexts from external knowledge bases.
@@
Real-world RAG systems often face significant challenges when handling large-scale, unstructured domain corpora (Peng et al. 2024; Zhang et al. 2025b; Su et al. 2025). Documents sourced from research papers, textbooks, or technical reports vary widely in reliability and completeness (Guo et al. 2025; Shen et al. 2025; Zhong et al. 2024; Wu et al. 2025), and the retrieved information is often complex and disorganized, as domain knowledge is typically scattered across multiple sources without clear dependencies (Sun et al. 2024; Ma et al. 2024; Hong et al. 2024). To manage this complexity, RAG systems commonly segment documents into smaller chunks for indexing (Borgeaud et al. 2022; Izacard et al. 2023; Jiang et al. 2023). This approach, however, sacrifices critical contextual information, leading to reduced retrieval accuracy and limited capability for complex reasoning tasks, particularly those requiring multi-hop reasoning across interconnected concepts.
To address this, recent advances (Zhang et al. 2024b; Procko and Ochoa 2024; Xiang et al. 2025; Bi et al. 2024c; Xiao et al. 2025a; Li et al. 2023a; Zhou et al. 2025b) leverage pre-constructed graphs to capture the relational connections among distributed documents, showing remarkable performance in complex tasks. The typical graph-based RAG (GraphRAG) approaches like Microsoft GraphRAG (Edge et al. 2024) utilize hierarchical community-based search and combine local and global querying to enhance response quality. Similarly, LightRAG (Guo et al. 2024) improves scalability through dual-level retrieval and graph-enhanced indexing, while HippoRAG (Gutiérrez et al. 2024) leverages PageRank-inspired algorithms to prioritize highly relevant nodes in memory-augmented retrieval for enhanced contextual coherence, RAPTOR (Sarthi et al. 2024) uses recursive abstractive processing for hierarchical text representation, and KGP (Wang et al. 2024) constructs domain-specific knowledge graphs with probabilistic reasoning to reduce noise. These GraphRAG methods excel in handling multi-hop queries by providing structured contextual depth.
@@
Despite recent advances, GraphRAG systems still face critical limitations in real-world scenarios. (i) Efficiency issues. Existing GraphRAG models rely on a costly process to transform the corpus into a graph, introducing overwhelming token cost and update latency as shown in Figure 1. It is hard to generalize to practical scenarios where knowledge bases are large-scale or dynamically evolving  (Edge et al. 2024). (ii) Low quality of the pre-built graph. Existing methods leverage LLMs to automatically build the graph without any guidance, which may introduce irrelevant or redundant information, leading to inefficiencies in both retrieval and reasoning (Guo et al. 2024). (iii) Lack of flexibility. Real-world queries vary in type and complexity, requiring different logic structures for accurate reasoning (Peng et al. 2024). The pre-built graph may not align with these required structures, resulting in ineffective knowledge retrieval. These challenges highlight the need for a more adaptive and efficient approach.
@@
During the subquery generation step in Equation 4, we consider two strategies for iterating over subproblems: (1) sampling with replacement, where the current batch of subproblems $S_{(i)}$ is retained after generating the unified subquery $q^{\text{uni}}_{(i)}$, allowing the language model to decide whether to proceed to the next batch $S_{(i+1)}$; and (2) sampling without replacement, where $S_{(i)}$ is removed after generating $q^{\text{uni}}_{(i)}$, enforcing forward progression to $S_{(i+1)}$. The first strategy, commonly adopted in iterative frameworks (Fu et al. 2023; Yang et al. 2024a; Li, Li, and Nie 2021), is prone to stalling at intermediate stages, producing near-duplicate subqueries until the maximum number of retrieval iterations is reached.
@@
Datasets. We evaluate our method’s RAG capabilities using three multi-hop question-answering benchmarks: MuSiQue (answerable) (Trivedi et al. 2022), 2WikiMultiHopQA (abbreviated as 2WikiMQA) (Ho et al. 2020), and HotpotQA (Yang et al. 2018). To manage experimental costs, we follow HippoRAG (Gutiérrez et al. 2024) by extracting 1,000 questions from each dataset’s validation set. Additionally, we adopt the approach of IRCOT (Trivedi et al. 2023) and HippoRAG (Gutiérrez et al. 2024) to gather all candidate passages, including both supporting and distractor passages, from the selected questions to create a retrieval corpus for each dataset. Dataset details are presented in Table 2.
@@
Baselines. We compare against several baselines: (i) Zero-shot LLM Inference (llama3(8b), llama3(13b) (LlamaIndex 2024), gpt3.5-turbo, gpt-4o-mini (Achiam et al. 2023)), (ii) vanilla retrieval augmented generation (top1, top3, top5), (iii) graph-based retrieval augmented generation (KGP (Wang et al. 2024), RAPTOR (Sarthi et al. 2024), G-retriever (He et al. 2024), GraphRAG (Edge et al. 2024), LightRAG (Guo et al. 2024), HippoRAG (Gutiérrez et al. 2024), HippoRAG2 (Gutiérrez et al. 2025)). For those who provide both single-step and multi-step options, we use their multi-step version for its general superior performance.
@@ B. Related Work
Graph-based RAG methods explicitly structure textual corpora into graphs to support multi-hop reasoning. For example, hierarchical-graph methods like GraphRAG (Edge et al. 2024) and RAPTOR (Sarthi et al. 2024) organize knowledge into multi-level document graphs (via community detection or recursive tree summarization) for coarse-to-fine retrieval. LightRAG (Guo et al. 2024) similarly builds a dual-level, graph-augmented index to enable efficient, incrementally updatable retrieval. Inspired by hippocampal memory models, HippoRAG (Gutiérrez et al. 2024) and HippoRAG2 (Gutiérrez et al. 2025) constructs a knowledge graph to index text chunks, and employ Personalized PageRank to perform one-shot multi-hop search. These graph-based approaches can employ off-the-shelf graph reasoning algorithms (Shengyuan et al. 2024; Chen et al. 2024b; Zhou et al. 2025a; Liu et al. 2023; Bei et al. 2023; Chen et al. 2024a; Li et al. 2023b) for enhancing retrieval accuracy. However, they rely on pre-built knowledge graphs or indexes. In contrast, LogicRAG eschews any fixed graph: it dynamically constructs a logic dependency graph of intermediate subqueries at inference time and uses a topological ordering to guide retrieval in a logically coherent sequence. LogicRAG also prunes redundant graph branches and context passages at each step to reduce overlap and token usage, enabling efficient multi-hop reasoning without a precomputed graph structure.
Reasoning-enhanced RAG methods instead interleave retrieval with the model’s own reasoning process. Think-on-Graph (Ma et al. 2024) alternates between using a static knowledge graph and text retrieval, leveraging entity links to deepen contextual retrieval iteratively. Likewise, RRP (Xiao et al. 2025c) distills guidance for LLM reasoning on well-established knowledge graphs. IRCoT (Trivedi et al. 2023) interleaves retrieval with a chain-of-thought: it uses the LLM’s intermediate reasoning steps to decide what to retrieve next, feeding retrieved facts back into the CoT. Likewise, IM-RAG (Yang et al. 2024b) trains the LLM to generate iterative “inner monologue” queries over multiple rounds of retrieval, refining its queries via reinforcement learning. Recent LAG (Xiao et al. 2025b) enhances problem solving by problem decomposition and logic augmentation from a Cartesian perspective. These methods improve multi-step QA by tightly coupling retrieval with reasoning steps, but they still depend on fixed resources (like a static KG) or predetermined retrieval schedules. LogicRAG, by contrast, constructs and updates its reasoning DAG adaptively during inference. It schedules subgoals by topologically sorting the DAG, effectively planning retrieval in dependency order, and prunes irrelevant branches and context on the fly. This dynamic, adaptive planning lets LogicRAG perform flexible multi-step reasoning efficiently, without any static graph or fixed retrieval plan.

### v2

@@ Abstract
Large language models (LLMs) often suffer from hallucination, generating factually incorrect statements when handling questions beyond their knowledge and perception. Retrieval-augmented generation (RAG) addresses this by retrieving query-relevant contexts from knowledge bases to support LLM reasoning. Recent advances leverage pre-constructed graphs to capture the relational connections among distributed documents, showing remarkable performance in complex tasks. However, existing Graph-based RAG (GraphRAG) methods rely on a costly process to transform the corpus into a graph, introducing overwhelming token cost and update latency. Moreover, real-world queries vary in type and complexity, requiring different logic structures for accurate reasoning. The pre-built graph may not align with these required structures, resulting in ineffective knowledge retrieval. To this end, we propose a Logic-aware Retrieval-Augmented Generation framework (LogicRAG) that dynamically extracts reasoning structures at inference time to guide adaptive retrieval without any pre-built graph. LogicRAG begins by decomposing the input query into a set of subproblems and constructing a directed acyclic graph to model the logical dependencies among them. To support coherent multi-step reasoning, LogicRAG then lin- earizes the graph using topological sort, so that subproblems can be addressed in a logically consistent order. Besides, LogicRAG applies graph pruning to reduce redundant retrieval and uses context pruning to filter irrelevant context, significantly reducing the overall token cost. Extensive experiments demonstrate that LogicRAG achieves both superior performance and efficiency compared to state-of-the-art baselines.
Codehttps://github.com/chensyCN/LogicRAG
Extended versionhttps://arxiv.org/abs/2508.06105
@@ Introduction
Large language models (LLMs), like Claude (anthropic2024claude) and ChatGPT (openai2023gpt4), have shown remarkable ability in a wide range of tasks like complex reasoning (LLM4EA; zhou2025self; zhou2025each; hong2024knowledge; hong2025next), question answering (khashabi2020unifiedqa), and social analysis (lu2021engage). However, these foundation models frequently struggle with domain-specific tasks, often producing hallucinations or inaccurate responses when handling knowledge-intensive queries (huang2023survey; fang2025safemlrm; fang2024alphaedit). To mitigate these issues, retrieval-augmented generation (RAG) (gao2023retrieval; lewis2020retrieval; zhang2025erarag) emerges as a promising framework that enhances LLMs by retrieving query-relevant contexts from external knowledge bases.
@@
Real-world RAG systems often face significant challenges when handling large-scale, unstructured domain corpora (peng2024graph; zhang2025faithfulrag). Documents sourced from research papers, textbooks, or technical reports vary widely in reliability and completeness (guo2025empowering; zhong2024unrealzoo; wu2025hierarchical), and the retrieved information is often complex and disorganized, as domain knowledge is typically scattered across multiple sources without clear dependencies (sun2024thinkongraph; ma2024think-on-graph-2.0). To manage this complexity, RAG systems commonly segment documents into smaller chunks for indexing (borgeaud2022improving; izacard2023atlas; jiang2023active). This approach, however, sacrifices critical contextual information, leading to reduced retrieval accuracy and limited capability for complex reasoning tasks, particularly those requiring multi-hop reasoning across interconnected concepts.
To address this, recent advances (zhang2024knowgpt; xiang2025use; li2023survey; zhou2025depth) leverage pre-constructed graphs to capture the relational connections among distributed documents, showing remarkable performance in complex tasks. The typical graph-based RAG (GraphRAG) approaches like Microsoft GraphRAG (edge2024local) utilize hierarchical community-based search and combine local and global querying to enhance response quality. Similarly HippoRAG (HippoRAG) leverages PageRank-inspired algorithms to prioritize highly relevant nodes in memory-augmented retrieval for enhanced contextual coherence, RAPTOR (sarthi2024raptor) uses recursive abstractive processing for hierarchical text representation, and KGP (wang2024knowledge) constructs domain-specific knowledge graphs with probabilistic reasoning to reduce noise. LinearRAG (linearrag) explores relation-free graph construction, eliminating LLM token costs and improve robust graphRAG with elegant bipartite graphs. These GraphRAG methods excel in handling multi-hop queries by providing structured contextual depth.
@@
Despite recent advances, GraphRAG systems still face critical limitations in real-world scenarios. (i) Efficiency issues. Existing GraphRAG models rely on a costly process to transform the corpus into a graph, introducing overwhelming token cost and update latency as shown in Figure 1. It is hard to generalize to practical scenarios where knowledge bases are large-scale or dynamically evolving  (edge2024local). (ii) Low quality of the pre-built graph. Existing methods leverage LLMs to automatically build the graph without any guidance, which may introduce irrelevant or redundant information, leading to inefficiencies in both retrieval and reasoning (guo2024lightrag). (iii) Lack of flexibility. Real-world queries vary in type and complexity, requiring different logic structures for accurate reasoning (peng2024graph). The pre-built graph may not align with these required structures, resulting in ineffective knowledge retrieval. These challenges highlight the need for a more adaptive and efficient approach.
@@
During the subquery generation step in Equation 4, we consider two strategies for iterating over subproblems: (1) sampling with replacement, where the current batch of subproblems $S_{(i)}$ is retained after generating the unified subquery $q^{\text{uni}}_{(i)}$, allowing the language model to decide whether to proceed to the next batch $S_{(i+1)}$; and (2) sampling without replacement, where $S_{(i)}$ is removed after generating $q^{\text{uni}}_{(i)}$, enforcing forward progression to $S_{(i+1)}$. The first strategy, commonly adopted in iterative frameworks (fu2023specializing; yang2024rag; li2021graph), is prone to stalling at intermediate stages, producing near-duplicate subqueries until the maximum number of retrieval iterations is reached.
@@
Datasets. We evaluate our method’s RAG capabilities using three multi-hop question-answering benchmarks: MuSiQue (answerable) (trivedi2022MuSiQue), 2WikiMultiHopQA (abbreviated as 2WikiMQA) (2wikimqa), and HotpotQA (yang2018hotpotqa). To manage experimental costs, we follow HippoRAG (HippoRAG) by extracting 1,000 questions from each dataset’s validation set. Additionally, we adopt the approach of IRCOT (IRCOT) and HippoRAG (HippoRAG) to gather all candidate passages, including both supporting and distractor passages, from the selected questions to create a retrieval corpus for each dataset. Dataset details are presented in Table 2.
@@
Baselines. We compare against several baselines: (i) Zero-shot LLM Inference (llama3(8b), llama3(13b) (llamaindex_website), gpt3.5-turbo, gpt-4o-mini (achiam2023gpt)), (ii) vanilla retrieval augmented generation (top1, top3, top5), (iii) graph-based retrieval augmented generation (KGP (wang2024knowledge), RAPTOR (sarthi2024raptor), G-retriever (he2024g), GraphRAG (edge2024local), LightRAG (guo2024lightrag), HippoRAG (HippoRAG), HippoRAG2 (hipporag2)). For those who provide both single-step and multi-step options, we use their multi-step version for its general superior performance.
@@ Conclusion
Acknowledgement
The work described in this paper was fully supported by a grant from the Innovation and Technology Commission of the Hong Kong Special Administrative Region, China (Project No. ITS/263/24FP).
@@ B. Related Work
Graph-based RAG methods explicitly structure textual corpora into graphs to support multi-hop reasoning. For example, hierarchical-graph methods like GraphRAG (edge2024local) and RAPTOR (sarthi2024raptor) organize knowledge into multi-level document graphs (via community detection or recursive tree summarization) for coarse-to-fine retrieval. LightRAG (guo2024lightrag) similarly builds a dual-level, graph-augmented index to enable efficient, incrementally updatable retrieval. Inspired by hippocampal memory models, HippoRAG (HippoRAG) and HippoRAG2 (hipporag2) constructs a knowledge graph to index text chunks, and employ Personalized PageRank to perform one-shot multi-hop search. These graph-based approaches can employ off-the-shelf graph reasoning algorithms (difflogic; neusymea; zhou2025taming; liu2023rsc; bei2023non; chen2024graphcrosscorelated; li2023gslb) for enhancing retrieval accuracy. However, they rely on pre-built knowledge graphs or indexes. In contrast, LogicRAG eschews any fixed graph: it dynamically constructs a logic dependency graph of intermediate subqueries at inference time and uses a topological ordering to guide retrieval in a logically coherent sequence. LogicRAG also prunes redundant graph branches and context passages at each step to reduce overlap and token usage, enabling efficient multi-hop reasoning without a precomputed graph structure.
Reasoning-enhanced RAG methods instead interleave retrieval with the model’s own reasoning process. Think-on-Graph (ma2024think-on-graph-2.0) alternates between using a static knowledge graph and text retrieval, leveraging entity links to deepen contextual retrieval iteratively. Likewise, RRP (xiao2025reliablereasoningpathdistilling) distills guidance for LLM reasoning on well-established knowledge graphs. IRCoT (IRCOT) interleaves retrieval with a chain-of-thought: it uses the LLM’s intermediate reasoning steps to decide what to retrieve next, feeding retrieved facts back into the CoT. Likewise, IM-RAG (im-rag) trains the LLM to generate iterative “inner monologue” queries over multiple rounds of retrieval, refining its queries via reinforcement learning. Recent LAG (xiao2025laglogicaugmentedgenerationcartesian) enhances problem solving by problem decomposition and logic augmentation from a Cartesian perspective. These methods improve multi-step QA by tightly coupling retrieval with reasoning steps, but they still depend on fixed resources (like a static KG) or predetermined retrieval schedules. LogicRAG, by contrast, constructs and updates its reasoning DAG adaptively during inference. It schedules subgoals by topologically sorting the DAG, effectively planning retrieval in dependency order, and prunes irrelevant branches and context on the fly. This dynamic, adaptive planning lets LogicRAG perform flexible multi-step reasoning efficiently, without any static graph or fixed retrieval plan.